OpenStudy (anonymous):

3x-6x=8 4x-8y=12 multiply the 1st by -4 and the 2nd by 3 and add results. How do I do that?

OpenStudy (compassionate):

3x - 6y = 8 4x - 8y = 12 -4(3x - 6y) = -4(8) 3(4x - 8y) = 3(12)

OpenStudy (anonymous):

-12x-12y=-32 12x-24y=36 ------------- then I just add them together right?

OpenStudy (compassionate):

Yes. Ad the. Notice your 12x's will cancel.

OpenStudy (anonymous):

-12x-24y=-32 12x-24y= 36 +-------------- -48y=4 is my answer but my multiple choices are: a)2y=-2 b)6y=6 c)0=0 d)0=4 I think d) but where did I go wrong? A negative and a negative need to be added.

OpenStudy (compassionate):

You have -48y = 4 Now solve for y

OpenStudy (compassionate):

i. 3x-6x=8 4x-8y=12 ii -4(3x - 6y) = 8 3(4x - 8y) = 12 iii. -12x + 24y = -32 12x - 24y = 36 iv. Now we simply add 12x -12x + 24y - 24y = 36 - 32

OpenStudy (anonymous):

I see it this way. Thanks a lot for your help. Wish I would've found this sight yesterday. I have been working on this worksheet for three days and I'm only on #9. Let's see how goes the next few problems.

OpenStudy (compassionate):

Well, welcome to OpenStudy! Ask anything you want and feel free to tag me!

OpenStudy (anonymous):

Thanks!